# 3D Voxel Maps to 2D Occupancy Maps for Efficient Path Planning for Aerial and Ground Robots

### v1

@@ Abstract
This article introduces a novel method for converting 3D voxel maps, commonly utilized by robots for localization and navigation, into 2D occupancy maps that can be used for more computationally efficient large-scale navigation, both in the sense of computation time and memory usage. The main aim is to effectively integrate the distinct mapping advantages of 2D and 3D maps to enable efficient path planning for both unmanned aerial vehicles (UAVs) and unmanned ground vehicles (UGVs). The proposed method uses the free space representation in the UFOMap mapping solution to generate 2D occupancy maps with height and slope information. In the process of 3D to 2D map conversion, the proposed method conducts safety checks and eliminates free spaces in the map with dimensions (in the height axis) lower than the robot’s safety margins. This allows an aerial or ground robot to navigate safely, relying primarily on the 2D map generated by the method. Additionally, the method extracts height and slope data from the 3D voxel map. The slope data identifies areas too steep for a ground robot to traverse, marking them as occupied, thus enabling a more accurate representation of the terrain for ground robots. The height data is utilized to convert paths generated using the 2D map into paths in 3D space for both UAVs and UGVs. The effectiveness of the proposed method is evaluated in two different environments.
@@ I INTRODUCTION
In the robotics literature, there are very few solutions for effectively converting 3D voxel maps into 2D maps. The most prevalent technique involves projecting the voxel map downward onto a 2D plane, a process implemented in the widely-used OctoMap library within ROS (Robot Operating System) [6]. Another example is the hybrid height voxel mapper, HMAPs [7], which allows for the conversion from 3D to 2.5D and 2D maps using downward projection. However, this approach has limitations. Specifically, the projected part of the map cannot include any voxels from the ground or ceiling, as these would also be projected onto the 2D map. Similarly, any obstacle in the map that is not within the selected range will be omitted from the projection. Generally, this is not a problem in smaller, controlled environments, but in larger environments with multiple height levels or in cases where there is drift in the robot’s positioning along the vertical axis, this method becomes less effective.
@@
While there are not many solutions for converting voxel-based 3D maps into 2D occupancy maps, several methods suitable for point clouds do exist. In the work [8], down projection is utilized to convert a 3D point cloud into a 2D map after filtering out the portion of the point cloud that represents the floor. In the past, multiple studies have also explored the conversion of point clouds generated by Kinect cameras  [9, 10, 11] into 2D maps. However, such methods are unsuitable for handling point clouds containing ceilings or overhangs. The method in [12] generates a 2D map using keyframes from ORB-slam [13], while the solution in  [7, 14] uses 2.5D mapping, focusing on mapping only surfaces that a ground robot can traverse.
@@
The article’s primary contribution is a novel method for converting a 3D voxel map into a 2D occupancy map, along with a height map, which contains the height of the environment’s floor and ceiling. The proposed method generates two 2D maps: i) an obstacle-free map that is useful for computation friendly UAV path planning by taking into account the height of the floor and ceiling from the height map, and ii) a map for UGVs that incorporates wall and obstacle detection using slope estimation, which again enables more computation-friendly path planning for UGVs in comparison with 3D Voxel maps. The second contribution is the proposal of a method to convert 2D paths generated using the 2D occupancy maps into 3D paths for UAVs and UGVs using the height map. These methods have been implemented in ROS [^1] Source code is available at https://github.com/LTU-RAI/3D-Voxel-Map-to-2D-Occupancy-Map-Conversion-Using-Free-Space-Representation, utilizing the UFOMap Mapping framework [15]. However, the method is compatible with any voxel-based mapping solution that represents space using occupied, free, and unknown voxel cells. The framework is particularly beneficial for computationally inexpensive and low-memory UGV navigation in 3D environments, especially in large-scale navigation scenarios or multi-agent operations with limited bandwidth, where map sharing is necessary. The efficacy of the proposed method has been validated in two scenarios, using voxel maps from a cave and an indoor environment.
This article is structured as follows: Section II presents the problem formulation. Section III describes the proposed methodology for 3D to 2D map conversion. Section IV discusses the utility of the 3D to 2D map conversions. Section V demonstrates the implementation of the methodology for converting 2D paths into 3D paths using the height map generated by the map conversion process. Section VI presents the validation of the overall methodology on maps from two different environments. Section VII contains a discussion of the results. Finally, Section VIII presents conclusions, with a discussion and summary of the findings.
@@
II Problem Formulation
The primary goal of this article is to convert a voxel-based 3D map into a 2D occupancy map, enhanced with height and slope information. This information is vital for path planning in UAVs and UGVs. Additionally, we utilize the 2D occupancy map to generate 3D paths for both aerial and ground robots. This involves transforming paths planned on the output 2D maps, which are produced by the proposed method using a 2D planner, into 3D paths. The map conversions consider whether the robot is aerial or terrestrial. Specifically, for UGVs, slopes that are too steep for traversal will be marked as occupied on the map. The methodology proposed in the article for map conversion is primarily designed for 3D maps produced by UFOMaps[15] but is compatible and generalizable to any voxel-based mapping solution that categorizes space as occupied, free, and unknown. Additionally, the method assumes that the 3D map does not contain multiple levels, such as a map of a single floor of a building or a cave with no overlapping paths.
@@ II Problem Formulation
An overview of the proposed methodology for 3D Voxel map to 2D map conversion is presented in Figure 1. Using the 3D voxel map $\mathcal{V}$, four 2D grid-based maps are generated. Firstly, a height map $\mathcal{H}$ is created, where every cell of a 2D grid map contains height data of the navigable free space relative to the origin of the Voxel map. Secondly, a slope map $\mathcal{S}$ is generated from the height map by estimating the slope of the floor plane in the immediate neighborhood of every cell. Finally, two occupancy maps $\mathcal{M}_{A}$ and $\mathcal{M}_{G}$ are produced, one for the UAVs and the other for UGVs respectively, assigning values: $-1$ for unknown cells, and a value in the range $[0,1]$ for the occupancy probability, with 00 being free, and $1$ being occupied. All 2D maps are generated over a 2D grid $\mathcal{G}_{2D}$ with the same $x$ and $y$ size/extent and resolution $\mathcal{V}_{res}$ as the Voxel map being converted. Details about the individual steps in the methodology are presented next.
@@ III-A Extraction of Free Space and Height Map
Most robots have their LiDAR mounted vertically, often resulting in an incomplete view of the floor. This configuration leads to numerous gaps in the mapped floor area, as can be seen in Figures 4(e) and 4(f) and Figures 5(e) and 5(f). To address this issue, instead of using occupied voxels, the free space is utilized to identify the floor in $\mathcal{V}$.
@@ III-A1 Free and occupied space
To identify the floor, the map $\mathcal{V}$ is first represented as a union of free $(\mathcal{F})$ and occupied $(\mathcal{O})$ ranges on every cell $\{x_{n},y_{n}\}$ in a 2D grid, denoted as $\mathcal{G}_{2D}$. The free space in the 3D map is represented as =$\mathcal{F}=\{F_{x_{1},y_{1}},F_{x_{2},y_{2}},\dots,F_{x_{n},y_{n}}\}$, where $F_{x_{n},y_{n}}=\{f_{1},f_{2},\dots,f_{i}\}$, and each $f_{i}$ is a range containing free spaces at 2D grid location $\{x_{n},y_{n}\}$. Each range of free space $f_{i}\in F_{x_{n},y_{n}}$ contains two values, $f_{i}=\{\hat{f}_{i},\check{f}_{i}\}$, where $\hat{f}_{i}$ is the higher and $\check{f}_{i}$ is the lower bound of the range. Similarly, the occupied space is represented as $\mathcal{O}=\{O_{x_{1},y_{1}},O_{x_{2},y_{2}},\dots,O_{x_{n},y_{n}}\}$, where $O_{x_{n},y_{n}}=\{o_{1},o_{2},\dots,o_{i}\}$ is the set of occupied ranges $o_{i}=\{\hat{o}_{i},\check{o}_{i}\}$. The ranges in the set $F_{x_{i},y_{i}}$ and in the set $O_{x_{i},y_{i}}$ cannot overlap or be connected with another range in the same set.
@@ III-A2 Height map
Since a 2D path planning algorithm utilizing the 2D maps generated in this work cannot perform collision checks on the height axis (the $Z$ axis), the proposed method disregards any free range $f_{i}\in\mathcal{F}$ with a height less than the safety margin $R_{maxZ}$, which is dependent on the dimensions of the robot. In other words, an $f_{i}$ is removed if $||f_{i}||<R_{maxZ}$, where $||f_{i}||$ denotes the height of the range. Given the ranges $F_{x_{i},y_{i}}$ and $O_{x_{i},y_{i}}$, the height map =$\mathcal{H}=\{h_{x_{1},y_{1}},h_{x_{2},y_{2}},\dots,h_{x_{n},y_{n}}\}$ is built as a collection of height ranges $h_{i}$, one for each cell $\{x_{i},y_{i}\}$, where $h_{i}=\{\hat{h}_{x_{n},y_{n}},\check{h}_{x_{n},y_{n}}\}$. The bottom of the height range $\check{h}_{x_{n},y_{n}}$ contains the height of the floor and $\hat{h}_{x_{n},y_{n}}$ is the height of the ceiling with respect to the origin of the voxel map. The floor of a cell $\{x_{n},y_{n}\}$ in the 2D grid is assumed to be the bottom of the lowest $f_{i}\in F_{x_{n},y_{n}}$, with the corresponding height set to $\check{h}_{x_{n},y_{n}}$, while the top of the same range $f_{i}\in F_{x_{n},y_{n}}$ is assumed to be the ceiling, with the corresponding height set to $\hat{h}_{x_{n},y_{n}}$.
Figure 2: The greedy approach to floor estimation might, in some situations, result in wrongly classifying the floor surface. In this case, the robot $R$ cannot see behind a fence, resulting in the floor (marked by red lines) being identified to be in the air.
Remark 1
As stated in the problem formulation in Section II, the assumption is that there is only one level of navigable space in the map. In environments where there are multiple layers of free space in the height axis, only the lowest free space is used by the method to generate the 2D maps.
The method for free space extraction and height mapping works exceptionally well (as will be presented in Section VI); however, due to its greedy nature, it has drawbacks in some scenarios. Figure 2 shows one such scenario, where a high fence obstructs the robot’s vision. In such a situation, the method will assume that the bottom of the visible parts is the floor, resulting in a perceived connected ’roof’ on the height map connected to the fence. Similarly, places at the edges of the lidar range will be considered as the floor.
@@ III-B Derivation of the Slope Map through Least Squares
The slope map, denoted as $\mathcal{S}=\{s_{x_{1},y_{1}},\dots,s_{x_{n},y_{n}}\}$, is a 2D grid where each cell, represented as $s_{x_{i},y_{i}}$, contains an estimate of the slope of the floor in a neighborhood of the cell. The slope at cell $s_{x_{i},y_{i}}$ is estimated as the slope of the plane fitted using the height and position information of the cells in the height map $\mathcal{H}$ at the same position as $s_{x_{i},y_{i}}$. The cells $h_{x_{j},y_{j}}$ that satisfy the condition $\max(|x_{j}-x_{i}|,|y_{j}-y_{i}|)\leq 1$ are considered as neighbours of the cell $s_{x_{i},y_{i}}$. To fit a plane at the cell $s_{x_{i},y_{i}}$, the linear equation of the form
@@
is considered, where the parameters $a$, $b$ and $c$ are real numbers. The interest here is in finding the slope of a plane that is a best fit for the height and position of the cells neighboring the cell $s_{x_{i},y_{i}}$, that is
$\underbrace{\begin{bmatrix}x_{j_{1}}&y_{j_{1}}&1\\ x_{j_{2}}&y_{j_{2}}&1\\ \vdots&\vdots&\vdots\end{bmatrix}}_{A}\begin{bmatrix}a\\ b\\ c\end{bmatrix}=\underbrace{\begin{bmatrix}\check{h}_{x_{j_{1}},y_{j_{1}}}\\ \check{h}_{x_{j_{2}},y_{j_{2}}}\\ \vdots\end{bmatrix}.}_{B}\$ (2)
The parameters of the best-fit plane are found using the least squares method, given by:
@@
Given Equation (1), the slope at cell $s_{x_{i},y_{i}}$ is given by Equation (4), where $a$ is the slope in the $x$ axis and $b$ is the slope in the $y$ axis.
$s_{x_{n},y_{n}}=\sqrt{a_{ls}^{2}+b_{ls}^{2}}\text{.}$ (4)
@@ III-C Generation of 2D Occupancy Maps for UAVs and UGVs
Two occupancy maps =$\mathcal{M}=\{m_{x_{1},y_{1}},\dots,m_{x_{n},y_{n}}\}$ are constructed: one for UAVs, $\mathcal{M}_{A}$ and the other for UGVs, $\mathcal{M}_{G}$.
@@ III-C1 UAV map $\mathcal{M}_{A}$
The map generation begins by representing all cells of the 2D occupancy maps with free space identified in subsection III-A as free, i.e., $m_{x_{n},y_{n}}=0$, when $F_{x_{j},y_{j}}\neq\{\varnothing\}$. The remaining cells $m_{x_{j},y_{j}}$ that are not free, but have at least one neighboring cell that is free are examined using the following criteria to decide their occupancy probabilities:
$m_{x_{j},y_{j}}=\left\{\begin{matrix}O_{cc}(x_{j},y_{j})\;&\text{if}\;O_{cc}(x_{j},y_{j})<O_{min}\\ -1\;&\text{else}\end{matrix}\right.$ (5)
where $O_{cc}(x_{j},y_{j})$ is the occupancy value in the range 00 and $1$ for the examined map cell $m_{x_{j},y_{j}}$ and $O_{min}$ is the minimum allowed occupancy value. The function of $O_{cc}(x_{j},y_{j})$ is defined in Equation (6) and calculates the overlapping percentage between the occupied ranges of $O_{x_{j},y_{j}}$ and the neighboring free space in the height map $\mathcal{H}$.
$O_{cc}(x_{j},y_{j})=\max_{{x_{k},y_{k}}\in N(x_{j},y_{j})}\left(\frac{\sum_{o_{i}\in O_{x_{j},y_{j}}}k(h_{x_{k},y_{k}},o_{i})}{||h_{x_{k},y_{k}}||}\right)$ (6)
where the function $k(f_{x_{n},y_{n}},o_{x_{n},y_{n}})$ is defined as:
$\begin{matrix}k(h_{x_{n},y_{n}},o_{i})=\max(0,\min(\hat{h}_{x_{n},y_{n}},\hat{o}_{i})-\max(\check{h}_{x_{n},y_{n}},\check{o}_{i}))\end{matrix}$ (7)
The function $k(h_{x_{n},y_{n}},o_{i})$ returns the length of the overlap between the two ranges, and when there is no overlap, the function returns 00. If the examined cell has multiple neighboring free cells, only the free cell that results in the maximum occupancy is used. In Equation (6), $N(x_{1},y_{1})$ is a set of neighboring cells $\{x_{n},y_{n}\}$ with $m_{x_{n},y_{n}}=0$, i.e., the cells containing free navigable range that satisfy the conditions in the following equation:
$\begin{matrix}N(x_{j},y_{j})=\\ \{x_{n},y_{n}|\max(|x_{n}-x_{j}|,|y_{n}-y_{j}|)=1.m_{x_{n},y_{n}}=0\}.\end{matrix}$ (8)

### v2

# Voxel Map to Occupancy Map Conversion Using Free Space Projection for Efficient Map Representation for Aerial and Ground Robots

@@ Abstract
This article introduces a novel method for converting 3D voxel maps, commonly utilized by robots for localization and navigation, into 2D occupancy maps for both unmanned aerial vehicles (UAVs) and unmanned ground vehicles (UGVs). The generated 2D maps can be used for more efficient global navigation for both UAVs and UGVs, in enabling algorithms developed for 2D maps to be useful in 3D applications, and allowing for faster transfer of maps between multiple agents in bandwidth-limited scenarios. The proposed method uses the free space representation in the UFOMap mapping solution to generate 2D occupancy maps. During the 3D to 2D map conversion, the method conducts safety checks and eliminates free spaces in the map with dimensions (in the height axis) lower than the robot’s safety margins. This ensures that an aerial or ground robot can navigate safely, relying primarily on the 2D map generated by the method. Additionally, the method extracts the height of navigable free space and a local estimate of the slope of the floor from the 3D voxel map. The height data is utilized in converting paths generated using the 2D map into paths in 3D space for both UAVs and UGVs. The slope data identifies areas too steep for a ground robot to traverse, marking them as occupied, thus enabling a more accurate representation of the terrain for ground robots. The effectiveness of the proposed method in enabling computationally efficient navigation for both aerial and ground robots is validated in two different environments, over both static maps and in online implementation in an exploration mission. The methods proposed within this article have been implemented in the popular robotics framework ROS and are open-sourced. The code is available at: https://github.com/LTU-RAI/Map-Conversion-3D-Voxel-Map-to-2D-Occupancy-Map.
@@ I INTRODUCTION
In the robotics literature, there are very few solutions for effectively converting 3D voxel maps into 2D maps. The most prevalent technique involves projecting a part of the voxel map at a fixed height downward onto a 2D plane, a process implemented in the widely-used OctoMap library within ROS (Robot Operating System) [6]. Another example is the hybrid height voxel mapper, HMAPs [7], which allows for the conversion from 3D to 2.5D and 2D maps using the same downward projection technique. Such approaches have several limitations. Specifically, the projected part of the map do not include any voxels from the ground or ceiling, as these would also be projected onto the 2D map. Similarly, any obstacle in the map that is not within the selected range will be omitted from the projection. Generally, this is not a problem in smaller, controlled environments, but in larger environments with varying height of navigable space, with multiple height levels or in cases where there is drift in the robot’s positioning along the vertical axis, this method becomes less effective.
@@
While there are not many solutions for converting voxel-based 3D maps into 2D occupancy maps, several methods suitable for point clouds do exist. In the work [8], down projection is utilized to convert a 3D point cloud into a 2D map, after filtering out the portion of the point cloud that represents the floor. In the past, multiple studies have also explored the conversion of point clouds generated by Kinect cameras  [9, 10, 11] into 2D maps. However, such methods are unsuitable for handling point clouds containing ceilings or overhangs. The method in [12] generates a 2D map using keyframes from ORB-slam [13], while the solution in  [7, 14] uses 2.5D mapping, focusing on mapping only surfaces that a ground robot can traverse.
II Contributions
@@
The article’s primary contribution is a novel method for converting a 3D voxel map into a 2D occupancy map. Instead of projecting a fixed section of the voxel map, as in existing methods, the proposed method projects free space, leading to a more robust and flexible approach in environments with varying elevations. The method also extracts the height values of the ceiling and floor of the environment and a local estimate of the slope of the floor. The proposed method generates two 2D maps: i) an obstacle-free map that is useful for computation-friendly UAV path planning, by taking into account the height of the floor and ceiling from the height map, and ii) a map for UGVs that incorporates walls and obstacles using slope estimation, which again enables more computation-friendly path planning for UGVs in comparison with 3D Voxel maps. The second contribution is the proposal of a method to convert 2D paths generated using 2D occupancy maps into 3D paths for UAVs and UGVs using the height map. This allows an aerial or ground robot to navigate safely, relying primarily on the 2D map generated by the method. These methods have been implemented in ROS and open sourced [^1] Source code is available at https://github.com/LTU-RAI/Map-Conversion-3D-Voxel-Map-to-2D-Occupancy-Map. The proposed method is implemented using the UFOMap Mapping framework [15]. However, the method is compatible with any voxel-based mapping solution that represents space using occupied, free, and unknown voxel cells.
@@
The rest of the article is structured as follows. Section III describes the proposed methodology for 3D to 2D map conversion. Section IV demonstrates the implementation of the methodology for converting 2D paths into 3D paths using the height map generated by the map conversion process. Section V discusses the utility of the 3D to 2D map conversions in robotics applications. Section VI presents the validation of the overall methodology on maps from two different environments. Section VII contains a discussion of the results. Finally, Section VIII presents conclusions, with a discussion and summary of the findings.
@@
An overview of the proposed methodology for 3D Voxel map to 2D map conversion is presented in Figure 1. Using the 3D voxel map $\mathcal{V}$, four 2D grid-based maps are generated. Firstly, a height map $\mathcal{H}$ is created, where every cell of a 2D grid map contains height data of the navigable free space (ceiling and floor) relative to the origin of the Voxel map. Secondly, a slope map $\mathcal{S}$ is generated from the height map (floor height) by estimating the slope of the floor plane in the immediate neighborhood of every cell. Finally, two occupancy maps $\mathcal{M}_{A}$ and $\mathcal{M}_{G}$ are produced, one for the UAVs and the other for UGVs respectively, assigning values: $-1$ for unknown cells, and a value in the range $[0,1]$ for the occupancy probability, with 00 being free, and $1$ being occupied. All 2D maps are generated over a 2D grid $\mathcal{G}_{2D}$ with the same $x$ and $y$ size/extent and resolution $\mathcal{V}_{res}$ as the Voxel map being converted. Details about the individual steps in the methodology are presented next.
@@ III-A Extraction of Free Space and Height Map
Most robots have their LiDAR mounted vertically, resulting in an incomplete view of the floor. This configuration leads to numerous gaps in the mapped floor area, as can be seen in Figures 4(e) and 4(f) and Figures 5(e) and 5(f). To address this issue, instead of using occupied voxels, the free space in the map is utilized to identify the floor in $\mathcal{V}$.
@@ III-A1 Free and occupied space
To identify the floor, the map $\mathcal{V}$ is first represented as a union of free $(\mathcal{F})$ and occupied $(\mathcal{O})$ ranges on every cell $\{x_{m},y_{n}\}$ in the 2D grid $\mathcal{G}_{2D}$ of size $M\times N$. The free space in the 3D map is represented by the set =$\mathcal{F}=\{F_{x_{m},y_{n}}\}_{M\times N}$, with the element $F_{x_{m},y_{n}}=\{f_{1},f_{2},\dots,f_{k_{mn}}\}$, where each $f_{i}\in F_{x_{m},y_{n}}$ is a range of free space along the height axis, at 2D grid location $\{x_{m},y_{n}\}$ and $k_{mn}$ is the number of such free space ranges. Each range of free space $f_{i}$ contains two values, $f_{i}=\{\hat{f}_{i},\check{f}_{i}\}$, where $\hat{f}_{i}$ and $\check{f}_{i}$ are the higher and lower bounds of the height range respectively. Similarly, the occupied space is represented as $\mathcal{O}=\{O_{x_{m},y_{n}}\}_{M\times N}$ where $O_{x_{m},y_{n}}=\{o_{1},o_{2},\dots,o_{l_{ij}}\}$ is the set of occupied ranges $o_{i}=\{\hat{o}_{i},\check{o}_{i}\}$. The ranges in the set $F_{x_{m},y_{n}}$ and in the set $O_{x_{m},y_{n}}$ cannot overlap or be connected with another range in the same set. Since a 2D path planning algorithm utilizing the 2D maps generated in this work cannot perform collision checks on the height axis (the $Z$ axis), the proposed method disregards any free range $f_{i}\in\mathcal{F}$ with a height less than the safety margin $R_{maxZ}$, which is dependent on the dimensions of the robot. In other words, an $f_{i}$ is removed if $|f_{i}|=(\hat{f}_{i}-\check{f}_{i})<R_{maxZ}$, where $|f_{i}|$ denotes the height of the range.
@@ III-A2 Height map
Given the free space ranges in $\mathcal{F}$, the height map =$\mathcal{H}=\{h_{x_{m},y_{n}}\}_{M\times N}$ is built as a collection of height ranges $h_{x_{m},y_{n}}$, one for each cell $\{x_{m},y_{n}\}$. An element of the height map $h_{x_{m},y_{n}}=\{\hat{h}_{x_{m},y_{n}},\check{h}_{x_{m},y_{n}}\}$, where the bottom of the height range $\check{h}_{x_{m},y_{n}}$ is the height of the floor and $\hat{h}_{x_{m},y_{n}}$ is the height of the ceiling with respect to the origin of the voxel map. The floor $\check{h}_{x_{m},y_{n}}$ of a cell $h_{x_{m},y_{n}}$ is the bottom of $f_{1}\in F_{x_{m},y_{n}}$, and the ceiling $\hat{h}_{x_{m},y_{n}}$ is the top of $f_{k_{mn}}\in F_{x_{m},y_{n}}$. The above formulation assumes only one level of navigable space in the map in the height axis, as in a building with just one floor or a subterranean environment with no overlapping tunnels.
@@ III-B Derivation of the Slope Map through Least Squares
The slope map, denoted as $\mathcal{S}=\{s_{x_{m},y_{n}}\}_{M\times N}$ is a 2D grid where each cell $s_{x_{m},y_{n}}$ contains an estimate of the slope of the floor in a neighborhood of the cell $s_{x_{m},y_{n}}$, which is the slope of the plane fitted using the floor height $\check{h}_{x_{i},y_{j}}$ and position $\{x_{i},y_{j}\}$ of the cells neighboring the position $\{x_{m},y_{n}\}$ in the grid. The cells $h_{x_{i},y_{j}}$ that satisfy the condition $\max(|x_{m}-x_{i}|,|y_{n}-y_{j}|)\leq S_{A}$ are considered when calculating the slope of the cell $s_{x_{m},y_{n}}$. $S_{A}\in\mathbb{N}_{\geq 1}$ is a tunable parameter for the size of the neighborhood considered when calculating the slope. To fit a plane at the cell $s_{x_{m},y_{n}}$, the linear equation of the form
@@
is considered, where the parameters $a$, $b$ and $c$ are real numbers. The interest here is in finding the slope of a plane that is a best fit for the height and position of the cells neighboring the cell $s_{x_{m},y_{n}}$, that is
$\underbrace{\begin{bmatrix}x_{i_{1}}&y_{j_{1}}&1\\ x_{i_{2}}&y_{j_{2}}&1\\ \vdots&\vdots&\vdots\end{bmatrix}}_{A}\begin{bmatrix}a\\ b\\ c\end{bmatrix}=\underbrace{\begin{bmatrix}\check{h}_{x_{i_{1}},y_{j_{1}}}\\ \check{h}_{x_{i_{2}},y_{j_{2}}}\\ \vdots\end{bmatrix}.}_{B}\$ (2)
The parameters of the best-fit plane are found using the least squares method, given by
@@
Given Equation (1), the slope at cell $s_{x_{m},y_{n}}$ is given by
$s_{x_{m},y_{n}}=\sqrt{a_{ls}^{2}+b_{ls}^{2}}\text{,}$ (4)
where $a_{ls}$ is the slope along the $x$ axis and $b_{ls}$ is the slope along the $y$ axis.
@@ III-C Generation of 2D Occupancy Maps for UAVs and UGVs
In this subsection, two occupancy maps =$\mathcal{M}=\{m_{x_{i},y_{j}}\}_{M\times N}$ are constructed: one for UAVs ($\mathcal{M}_{A}$) and the other for UGVs ($\mathcal{M}_{G}$).
@@ III-C1 UAV map $\mathcal{M}_{A}$
The map generation begins by representing all cells of the 2D occupancy map with free space identified in subsection III-A as free, i.e., $m_{x_{i},y_{j}}=0$, when $F_{x_{i},y_{j}}\neq\{\varnothing\}$ i.e, when the set $F_{x_{i},y_{j}}$ is non-empty. The remaining cells $m_{x_{i},y_{j}}$ that are not free, but have at least one neighboring cell that is free are examined using the following criteria to decide their occupancy probabilities:
$m_{x_{i},y_{j}}=\left\{\begin{matrix}O_{cc}(x_{i},y_{j})\;&\text{if}\;O_{cc}(x_{i},y_{j})\geq O_{min}\\ -1\;&\text{otherwise}\end{matrix}\right.$ (5)
where $O_{cc}(x_{i},y_{j})$ is the occupancy value in the range 00 and $1$ for the examined map cell $m_{x_{i},y_{j}}$ computed using Equation (6) and $O_{min}$ is the minimum allowed occupancy value. If $O_{cc}(x_{i},y_{j})<O_{min}$, the cell’s occupancy is deemed unknown and the value -1 is assigned. The function $O_{cc}(x_{i},y_{j})$, which calculates the overlapping percentage between the occupied ranges of $O_{x_{i},y_{j}}$ and the neighboring free space in the height map $\mathcal{H}$, is defined as
$O_{cc}(x_{i},y_{j})=\max_{{x_{m},y_{n}}\in N(x_{i},y_{j})}\left(\frac{\sum_{o_{i}\in O_{x_{i},y_{j}}}k(h_{x_{m},y_{n}},o_{i})}{||h_{x_{m},y_{n}}||}\right),$ (6)
where the function $k(h_{x_{m},y_{n}},o_{i})$ defined as:
$\begin{matrix}k(h_{x_{m},y_{n}},o_{i})=\max(0,\min(\hat{h}_{x_{m},y_{n}},\hat{o}_{i})-\max(\check{h}_{x_{m},y_{n}},\check{o}_{i}))\end{matrix}$ (7)
returns the length of the overlap between the two ranges, and when there is no overlap, the function returns 00. If the examined cell has multiple neighboring free cells, only the free cell that results in the maximum occupancy is used. In Equation (6), $N(x_{i},y_{j})$ is a set of neighboring cells $\{x_{r},y_{s}\}$ with $m_{x_{r},y_{s}}=0$, i.e., the cells containing free navigable range that satisfy the conditions in the following equation:
$N(x_{i},y_{j})=$ (8)
$\{\{x_{r},y_{s}\}|\max$ $(|x_{r}-x_{i}|,|y_{s}-y_{j}|)=1,m_{x_{r},y_{s}}=0\}.$